# Alternating Minimization Schemes for Computing Rate-Distortion-Perception Functions with $f$-Divergence Perception Constraints

### v1

@@ II-C Alternating Minimization and BA-type algorithms
where $\mathcal{X}$ and $\mathcal{Y}$ are two arbitrary non-empty sets and the function $f(x,y)$ satisfies $-\infty<f(x,y)\leq+\infty$ for each $x\in\mathcal{X}$ and $y\in\mathcal{Y}$. Furthermore, we assume that, for each $x\in\mathcal{X}$, there exists $y\in\mathcal{Y}$ with $f(x,y)<+\infty$, implying that $s\coloneqq\displaystyle\inf_{\begin{subarray}{c}x\in\mathcal{X},y\in\mathcal{Y}\end{subarray}}f(x,y)<+\infty$. Depending on the case, the existence or uniqueness of the minimizer $(x^{*},y^{*})$ such that $f(x^{*},y^{*})=s$ may also be assumed.
The goal of the alternating minimization method is to create a sequence $\{(x^{(n)},y^{(n)})\}$ such that $\lim_{n\to\infty}f(x^{(n)},y^{(n)})=s$. Under specific conditions, such sequence can be defined using the solutions of two sub-problems; for $x_{i}\in\mathcal{X}$, $h(x_{i})=\displaystyle\operatorname*{arg\,min}_{y\in\mathcal{Y}}f(x_{i},y)$ and, for $y_{i}\in\mathcal{Y}$, $g(y_{i})=\displaystyle\operatorname*{arg\,min}_{x\in\mathcal{X}}f(x,y_{i})$. For some initial point $y^{(0)}$, we can define the $n$-th element of the sequence as:
$x^{(n)}=g(y^{(n-1)})\qquad y^{(n)}=g(x^{(n)})\qquad\text{for }n=1,2\ldots.$
@@
  • In [37], Csiszár and Tusnády prove that, if the sequence $\{(x^{(n)},y^{(n)})\}$ guarantees
    $f(x,y)+f(x,y^{(n-1)})\geq f(x,y^{(n)})+f(x^{(n)},y^{(n-1)})\quad\forall x\in\mathcal{X},\forall y\in\mathcal{Y}$
@@
  • In [38], Grippo and Sciandrone prove convergence of the sequence to a stationary point of $f$, under the assumption of convexity of the feasible sets $\mathcal{X}$ and $\mathcal{Y}$ and existence of the sequence limit.
BA-type algorithms, introduced for the numerical computation of channel capacity curve [39] and RD curve [22], represent specific instances of alternating minimization algorithms [40, Chapter 9]. In fact, in their classic formulation, both problems can be expressed as constrained minimization of a convex function on the sets of marginal distributions and transition matrices, where the properties of the sets (e.g., convexity) depend on the constraints for which the problem is formulated.
@@ III Main Results
In this section, we present the derivation of our main theoretical results. We start by providing in the following lemma the parametric characterization of the solution of the RDPF problem, obtained by casting (1) as a double minimization problem.
@@
(Double minimization) Let $D\geq 0$, $P\geq 0$ and let $D(\cdot||\cdot)=D_{f}(\cdot||\cdot)$. Moreover, let $s=(s_{D},s_{P})$ with $s_{D}\geq{0}$, $s_{P}\geq{0}$ being the Lagrangian multipliers associated with constraints (2) and (3). Then (1) can be expressed as a double minimum as
$\begin{split}R(D,P)=\min_{Q_{Y|X}\in\mathcal{Q}(\mathcal{X})}\min_{h\in\mathcal{P}(\mathcal{X})}D_{KL}(p_{X}Q_{Y|X}||p_{X}h)&+s_{D}\left(\mathbb{E}\left[d(x,y)\right]-D\right)+s_{P}(D_{f}(p_{X}||q_{Y})-P)\end{split}$ (4)
where $D=\mathbb{E}_{Q_{Y|X}^{*}}\left[d(x,y)\right]$ and $P=D_{f}(p_{X}||q_{Y}^{*})$, with $Q_{Y|X}^{*}$ achieving the minimum.
Furthermore, for fixed $Q_{Y|X}$, the right side of (4) is minimized by
@@
while, for fixed $h$, the right side of (4) is minimized by
@@
We remark that, although showing a close resemblance to the classical BA solution [41, Theorem 6.3.3], Lemma 1 differs from it in (6). In particular, the perception constraint (3) induces the presence of an additional exponential term, i.e., $s_{P}g(\cdot,\cdot)$. Clearly, the classical BA implicit solution can be obtained as a special case of (6) by considering $s_{P}=0$, effectively removing the perceptual constraint. The next corollary follows as a direct consequence of Lemma 1.
@@
$R(D,P)$ in (4) can be reformulated as follows
$\begin{split}R(D_{s},P_{s})=-s_{D}D_{s}-s_{P}P_{s}+\min_{h\in\mathcal{P}(\mathcal{X})}&\quad s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\partial f\left(\frac{p_{X}(y)}{q_{Y}[h](y)}\right)\\ &-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}h(y)A[q_{Y}[h]](x,y,s)\right)\end{split}$ (9)
where
$D_{s}$ $=\sum_{(x,y)\in\mathcal{X}^{2}}\frac{p_{X}(x)h^{*}(y)A[h^{*}](x,y,s)}{\sum_{i\in\mathcal{X}}h^{*}(i)A[h^{*}](x,i,s)}d(x,y)\qquad P_{s}=D_{f}(p_{X}||h^{*})$
@@
The proof follows by substitution of (6) in (4). ∎
The following lemma characterizes a necessary and sufficient condition to ensure that, for given Lagrangian multipliers $s=(s_{D},s_{P})$, a distribution $h^{*}\in\mathcal{P}(\mathcal{X})$ is the optimal solution of (9), i.e., $(h^{*},Q_{Y|X}[h^{*}])$ defines a point achieving the RDPF.
@@
$\frac{h^{(n+1)}(y)}{h^{(n)}(y)}=\sum_{x\in\mathcal{X}}\frac{p_{X}(x)\exp\left\{-s_{D}d(x,y)-s_{P}g\left(p_{X},{h^{(n+1)}},y\right)\right\}}{\sum_{i\in\mathcal{X}}h^{(n)}(i)\exp\left\{-s_{D}d(x,i)-s_{P}g\left(p_{X},{h^{(n+1)}},i\right)\right\}}=c[h^{(n)},h^{(n+1)}](y)$ (11)
@@ III-A NAM scheme
The implicit definition of $h^{(n+1)}$ in (11) suggests the application of numerical methods for its approximation. To this end, we introduce a variation to the OAM scheme, referred to as NAM scheme, where $h^{(n+1)}$ is approximated at each minimization step using Newton’s root-finding method [32].
@@
The application of Newton’s method requires the existence and the invertibility of the Jacobian matrix $J_{T}$ of the functional $T$ [32, Section 10.2]. In our case, ensuring the existence of $J_{T}$ requires a more restrictive continuity assumption on the divergence measure, i.e., $D_{f}(\cdot||\cdot)$ needs to be twice differentiable in its second argument. Although this limitation reduces the generality of the NAM scheme, we note that the majority of the commonly used divergences, i.e., see Section II-B, respect this assumption. Under this restriction, the invertibility of $J_{T}$ is shown in the following lemma.
@@
$M[h^{(n)},u]$
@@
$u^{(k+1)}\triangleq u^{(k)}-\left(J_{T}[h^{(n)},u^{k}]\right)^{-1}\cdot T[h^{(n)},u^{(k)}].$
@@
The proof follows by direct application of Newton’s root-finding method [32, Section 10.2], since Lemma 3 proves that the set of solutions $h^{(n+1)}$ and the set of the roots of $T(\cdot)$ coincide, while Lemma 4 proves that $T(\cdot)$ satisfies the assumption for the convergence of the method. ∎
@@ III-B RAM scheme
where $A[\cdot]$ is defined in (7) and $v[\cdot]:\mathcal{P}(\mathcal{X})\to\mathcal{P}(\mathcal{X})$ is any functional defining a probability distribution. Let $\hat{h}^{(0)}$ be any probability vector with nonzero components and let $\hat{Q}_{Y|X}^{(n+1)}\equiv\hat{Q}_{Y|X}[\hat{h}^{(n)}]$, $\hat{h}^{(n+1)}\equiv\hat{q}_{Y}[\hat{h}^{(n)}]$, and $v^{(n)}=v[h^{(n)}]$. Then, as $n\xrightarrow{}\infty$, we obtain
@@
if $\lim_{n\rightarrow\infty}||h^{(n+1)}-v^{(n)}||=0$ with at least linear rate of convergence.
@@
Theorem 4 enables the implementation of the alternating minimization scheme by introducing an auxiliary variable $v[h^{(n)}]$, which approximates the correct iteration $h^{(n+1)}$ while still being a function of only the current iteration of $h^{(n)}$. Nevertheless, depending on $v$, this approximation may incur restrictions on the domain of the Lagrangian multiplier $s_{P}$ that affect convergence guarantees, as we will later discuss in Section IV.
@@
(NAM vs. RAM) The main advantage of NAM is the convergence for any value of the Lagrangian multipliers $(s_{D},s_{P})$, without needing any additional condition. However, the introduction of Newton’s method requires the differentiability of the perceptual metric and the introduction of additional complexity at each iteration. RAM, on the other hand, lifts the differentiability requirement and avoids the additional computational cost at each iteration, but at the expense of a potentially smaller set of $(s_{D},s_{P})$ for which the algorithm achieves the optimal solution, potentially prohibiting the computation of the complete RDP curve.
@@ IV Algorithmic Implementation and Convergence Analysis
This section is about the algorithmic implementation of the alternating minimization schemes derived in Section III and the characterization of their convergence rate.
@@
1:source distribution $p_{X}$; Lagrangian parameters $s=(s_{D},s_{P})$ with $s_{D}\geq 0$ and $s_{P}\geq 0$; error tolerances $\epsilon>0$, $\delta>0$; distortion measure $d(x,y)$; initial assignment $h^{(0)}$.
2:$\omega\leftarrow+\infty$; $n\leftarrow 0$;
3:while $\omega\geq\epsilon$ do
4:     $h^{(n+1)}\leftarrow$ Newton Approximation( $p_{X}$, $h^{(n)}$, $s$, $\delta$)
5:     $c^{(n)}\leftarrow c\left[h^{(n)},h^{(n+1)}\right]$
6:     $\omega\leftarrow\log c^{(n)}_{\max}(y)-\sum_{y\in\mathcal{X}}h^{(n)}c^{(n)}(y)\log(c^{(n)}(y))$
7:     $n\leftarrow n+1$
8:end while
@@
2:$\omega\leftarrow+\infty$; $n\leftarrow 0$;
3:while $\omega\geq\epsilon$ do
4:     $c^{(n)}\leftarrow c[\hat{h}^{(n)},v^{(n)}]$
5:     $\hat{h}^{(n+1)}\leftarrow\hat{h}^{(n)}\cdot c^{(n)}$
6:     $\omega\leftarrow\log c^{(n)}_{\max}(y)-\sum_{y\in\mathcal{X}}\hat{h}^{(n)}c^{(n)}(y)\log(c^{(n)}(y))$
7:     $n\leftarrow n+1$
8:end while
@@
We first focus on the derivation of a stopping criterion for the RAM case, as the NAM case can be obtained by fixing the auxiliary variable $v[h^{(n)}]=q_{Y}[h^{(n)}]$ in Theorem 4, i.e., recovering the original OAM iterates. For this purpose, we need the following theorem in which we derive bounds on the RDPF.
@@
(Bounds on RDPF) Let $\hat{Q}_{Y|X}$ and $\hat{h}$ be defined as in Theorem 4, $c()$ be as defined as in Lemma 2, and $c_{\max}=\max_{y\in\mathcal{X}}c[u,\hat{q}_{Y}[u]](y)$. Then, at the point $D=\mathbb{E}_{\hat{Q}_{Y|X}}\left[d(x,y)\right]$, and $P=D_{f}(p_{X}||\hat{q}_{Y})$, the following bounds hold
$R(D,P)$ $\geq R_{L}[u](D,P)=-s_{D}D-s_{P}P+\hat{W}[u]-\log(c_{\max})$ (17)
$\begin{split}R(D,P)&\leq R_{U}[u](D,P)=-s_{D}D-s_{P}P+\hat{W}[u]\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\sum_{y\in\mathcal{X}}u(y)c[u,\hat{q}_{Y}[u]](y)\log(c[u,\hat{q}_{Y}[u]](y))\end{split}$ (18)
where $\hat{W}{}[\hat{h}]$ is given by
$\begin{split}\hat{W}[u]&=-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}u(y)A^{(n)}(x,,s)\right)+s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\frac{p_{X}(y)}{u(y)}\partial f\left(\frac{p_{X}(y)}{u(y)}\right)\\ &\quad+s_{P}\Bigg{[}\sum_{y\in\mathcal{X}}\hat{q}_{Y}[u]\left(f\left(\frac{p_{X}(y)}{q_{Y}[u](y)}\right)-f\left(\frac{p_{X}(y)}{u(y)}\right)\right)\Bigg{]}.\end{split}$ (19)

### v2

@@ II-C Alternating Minimization and BA-type algorithms
where $\mathcal{X}$ and $\mathcal{Y}$ are two arbitrary non-empty sets and the function $f(x,y)$ satisfies $\infty<f(x,y)\leq+\infty$ for each $x\in\mathcal{X}$ and $y\in\mathcal{Y}$. Furthermore, we assume that, for each $x\in\mathcal{X}$, there exists $y\in\mathcal{Y}$ with $f(x,y)<+\infty$, implying that $s\coloneqq\displaystyle\inf_{\begin{subarray}{c}x\in\mathcal{X},y\in\mathcal{Y}\end{subarray}}f(x,y)<+\infty$. Depending on the case, the existence or uniqueness of the minimizer $(x^{*},y^{*})$ such that $f(x^{*},y^{*})=s$ may also be assumed.
The goal of the alternating minimization method is to construct a sequence $\{(x^{(n)},y^{(n)})\}$ such that $\lim_{n\to\infty}f(x^{(n)},y^{(n)})=s$. Under specific conditions, such a sequence can be defined using the solutions of two sub-problems; for $x_{i}\in\mathcal{X}$, $h(x_{i})=\displaystyle\operatorname*{arg\,min}_{y\in\mathcal{Y}}f(x_{i},y)$ and, for $y_{i}\in\mathcal{Y}$, $g(y_{i})=\displaystyle\operatorname*{arg\,min}_{x\in\mathcal{X}}f(x,y_{i})$. Starting from an initial point $y^{(0)}$, we can define the $n$-th element of the sequence as:
$x^{(n)}=g(y^{(n-1)})\qquad y^{(n)}=h(x^{(n)})\qquad\text{for }n=1,2\ldots.$
@@
  • In [39], Csiszár and Tusnády prove that, if the sequence $\{(x^{(n)},y^{(n)})\}$ guarantees $\forall x\in\mathcal{X},\forall y\in\mathcal{Y}$
    $f(x,y)+f(x,y^{(n-1)})\geq f(x,y^{(n)})+f(x^{(n)},y^{(n-1)}),$
@@
  • In [40], Grippo and Sciandrone prove the convergence of the sequence to a stationary point of $f$, under the assumption of convexity of the feasible sets $\mathcal{X}$ and $\mathcal{Y}$ and existence of the sequence limit.
BA-type algorithms, introduced for the numerical computation of channel capacity [41] and RD function [24], are specific instances of alternating minimization algorithms [42, Chapter 9]. In fact, in their classic formulation, both problems can be expressed as constrained minimization of a convex function on the sets of marginal distributions and transition matrices, where the properties of the sets (e.g., convexity) depend on the constraints for which the problem is formulated.
@@ III Main Results
In this section, we present the derivation of our main theoretical results. We start by providing the parametric characterization of the solution of the RDPF problem in the following lemma, obtained by casting (1) as a double minimization problem.
@@
(Double minimization) Let $D\geq 0$, $P\geq 0$ and let $D(\cdot||\cdot)=D_{f}(\cdot||\cdot)$. Moreover, let $s=(s_{D},s_{P})$ with $s_{D}\geq{0}$, $s_{P}\geq{0}$ being the Lagrangian multipliers associated with constraints (2) and (3). Then (1) can be expressed as a double minimum
$\begin{split}R(D,P)=\min_{\begin{subarray}{c}Q_{Y|X}\in\mathcal{Q}(\mathcal{X})\\ h\in\mathcal{P}(\mathcal{X})\end{subarray}}&D_{KL}(p_{X}Q_{Y|X}||p_{X}h)\\ &+s_{D}\left(\mathbb{E}\left[d(X,Y)\right]-D\right)\\ &+s_{P}(D_{f}(p_{X}||q_{Y})-P)\end{split}$ (4)
where $D=\mathbb{E}_{Q_{Y|X}^{*}}\left[d(X,Y)\right]$ and $P=D_{f}(p_{X}||h^{*})$, with $(Q_{Y|X}^{*},h^{*})$ being the pair achieving the minimum.
Furthermore, for fixed $Q_{Y|X}$, the right-hand side of (4) is minimized by
@@
whereas, for fixed $h$, the right-hand side of (4) is minimized by
@@
We remark that, although showing a close resemblance to the classical BA solution [43, Theorem 6.3.3], Lemma 1 differs from it in (6). In particular, the perception constraint (3) induces the presence of an additional exponential term, i.e., $s_{P}g(\cdot,\cdot)$. Note that the classical BA implicit solution can be obtained as a special case of (6) by considering $s_{P}=0$, effectively removing the perceptual constraint. The next corollary follows as a direct consequence of Lemma 1.
@@
Let $s=(s_{D},s_{P})$ with $s_{D}\geq{0}$, $s_{P}\geq{0}$. Then $R(D,P)$ in (4) can be reformulated as follows
$R(D_{s},P_{s})=$ $-s_{D}D_{s}-s_{P}P_{s}$ (9)
$+\min_{h\in\mathcal{P}(\mathcal{X})}$ $\quad s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\partial f\left(\frac{p_{X}(y)}{q_{Y}[h](y)}\right)$
$-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}h(y)A[q_{Y}[h]](x,y,s)\right)$
where $P_{s}=D_{f}(p_{X}||h^{*})$ and
$D_{s}$ $=\sum_{(x,y)\in\mathcal{X}^{2}}\frac{p_{X}(x)h^{*}(y)A[h^{*}](x,y,s)}{\sum_{i\in\mathcal{X}}h^{*}(i)A[h^{*}](x,i,s)}d(x,y)$
@@
The proof follows by substitution of (6) into (4). ∎
We note that in Corollary 1 and in subsequent analysis, the subscript notation $(D_{s},P_{s})$ is introduced to explicitly indicate the dependence of the constraint levels $(D,P)$ on the fixed Lagrangian multipliers $s=(s_{D},s_{P})$. The following lemma characterizes a necessary and sufficient condition to ensure that, for given Lagrangian multipliers $s=(s_{D},s_{P})$, a distribution $h^{*}\in\mathcal{P}(\mathcal{X})$ is the optimal solution of (9), i.e., $(h^{*},Q_{Y|X}[h^{*}])$ defines a point achieving the RDPF.
@@
$\frac{h^{(n+1)}(y)}{h^{(n)}(y)}$ $=\sum_{x\in\mathcal{X}}\frac{p_{X}(x)e^{-s_{D}d(x,y)-s_{P}g\left(p_{X},{h^{(n+1)}},y\right)}}{\sum_{i\in\mathcal{X}}h^{(n)}(i)e^{-s_{D}d(x,i)-s_{P}g\left(p_{X},{h^{(n+1)}},i\right)}}$
$=c[h^{(n)},h^{(n+1)}](y)$ (11)
@@ III-A NAM scheme
The implicit definition of $h^{(n+1)}$ in (11) suggests the application of numerical methods for its approximation. To this end, we introduce a variation of the OAM scheme, referred to as NAM scheme, where $h^{(n+1)}$ is approximated at each minimization step using Newton’s root-finding method [33].
@@
The application of Newton’s method requires the existence and the invertibility of the Jacobian matrix $J_{T}$ of the functional $T$ [33, Section 10.2]. In our case, ensuring the existence of $J_{T}$ requires a more restrictive continuity assumption on the divergence measure, i.e., $D_{f}(\cdot||\cdot)$ needs to be twice differentiable in its second argument. Although this limitation reduces the generality of the NAM scheme, we note that most commonly used divergences (see Section II-B) satisfy this assumption. Under this restriction, the invertibility of $J_{T}$ is shown in the following lemma.
@@
$M[h^{(n)},u]=$ (14)
$~\left[h^{(n)}(i)\sum_{x\in\mathcal{X}}p_{X}(x)\frac{A[u](x,i,s)\cdot A[u](x,j,s)}{\left(\sum_{k\in\mathcal{X}}h^{(n)}(k)A[u](x,k,s)\right)^{2}}\right]$
@@
$u^{(k+1)}\triangleq u^{(k)}-\left(J_{T}[h^{(n)},u^{(k)}]\right)^{-1}\cdot T[h^{(n)},u^{(k)}].$
@@
The proof follows by direct application of Newton’s root-finding method [33, Section 10.2], since Lemma 3 proves that the set of solutions $h^{(n+1)}$ and the set of the roots of $T(\cdot)$ coincide, while Lemma 4 proves that $T(\cdot)$ satisfies the assumption for the convergence of the method. ∎
@@ III-B RAM scheme
where $A[\cdot]$ is defined in (7) and $v[\cdot]:\mathcal{P}(\mathcal{X})\to\mathcal{P}(\mathcal{X})$ is any functional defining a probability distribution. Let $\hat{h}^{(0)}$ be any probability vector with nonzero components and let $\hat{Q}_{Y|X}^{(n+1)}\equiv\hat{Q}_{Y|X}[\hat{h}^{(n)}]$, $\hat{h}^{(n+1)}\equiv\hat{q}_{Y}[\hat{h}^{(n)}]$, and $v^{(n)}=v[\hat{h}^{(n)}]$. Then, as $n\xrightarrow{}\infty$, we obtain
@@
[…]
@@
Theorem 4 enables the implementation of the alternating minimization scheme by introducing an auxiliary variable $v[h^{(n)}]$, which approximates the correct iteration $h^{(n+1)}$ while still being a function of only the current iteration of $h^{(n)}$. Nevertheless, depending on $v$, this approximation may incur restrictions on the domain of the Lagrangian multiplier $s_{P}$ that affect convergence guarantees, as discussed later in Section IV.
@@
(NAM vs. RAM) The main advantage of NAM is the convergence for any value of the Lagrangian multipliers $(s_{D},s_{P})$, without the need for any additional condition. However, the introduction of Newton’s method requires the differentiability of the perceptual metric and the introduction of additional complexity at each iteration. RAM, on the other hand, removes the differentiability requirement and avoids the additional computational cost at each iteration, but at the expense of a potentially smaller set of $(s_{D},s_{P})$ for which the algorithm achieves the optimal solution, which may preclude the computation of the complete RDP curve.
@@ IV Algorithmic Implementation and Convergence Analysis
This section addresses the algorithmic implementation of the alternating minimization schemes derived in Section III and the characterization of their convergence rate.
@@
1:source distribution $p_{X}$; Lagrangian parameters $s=(s_{D},s_{P})$ with $s_{D}\geq 0$ and $s_{P}\geq 0$; error tolerances $\epsilon>0$, $\delta>0$; distortion measure $d(\cdot,\cdot)$; initial assignment $h^{(0)}$.
2:
[…]
4:while $\omega\geq\epsilon$ do
5:  $h^{(n+1)}\leftarrow$ Newton Approx.( $p_{X}$, $h^{(n)}$, $s$, $\delta$)
6:  $c^{(n)}\leftarrow c\left[h^{(n)},h^{(n+1)}\right]$
7:  $\omega\leftarrow\log c^{(n)}_{\max}(y)-\sum_{y\in\mathcal{X}}h^{(n)}c^{(n)}(y)\log(c^{(n)}(y))$
8:  $n\leftarrow n+1$
9:end while
10:
@@
2:
[…]
4:while $\omega\geq\epsilon$ do
5:  $c^{(n)}\leftarrow c[\hat{h}^{(n)},v^{(n)}]$
6:  $\hat{h}^{(n+1)}\leftarrow\hat{h}^{(n)}\cdot c^{(n)}$
7:  $\omega\leftarrow\log c^{(n)}_{\max}(y)-\sum_{y\in\mathcal{X}}\hat{h}^{(n)}c^{(n)}(y)\log(c^{(n)}(y))$
8:  $n\leftarrow n+1$
9:end while
10:
@@
We first derive a stopping criterion for the RAM case, since the NAM case can be obtained by fixing the auxiliary variable $v[h^{(n)}]=q_{Y}[h^{(n)}]$ in Theorem 4, i.e., recovering the original OAM iterates. For this purpose, we need the following theorem which establishes bounds on the RDPF.
@@
[…]
$\begin{split}R(D,P)&\geq R_{L}[\hat{h}](D,P)\\ &=-s_{D}D-s_{P}P+\hat{W}[\hat{h}]-\log(c_{\max})\end{split}$ (18)
$\begin{split}R(D,P)&\leq R_{U}[\hat{h}](D,P)\\ &=-s_{D}D-s_{P}P+\hat{W}[\hat{h}]\\ &~~~~-\sum_{y\in\mathcal{X}}\hat{h}(y)c[\hat{h},\hat{q}_{Y}[\hat{h}]](y)\log(c[\hat{h},\hat{q}_{Y}[\hat{h}]](y))\end{split}$ (19)
where $\hat{W}{}$$\cdot$] is given by
$\begin{split}\hat{W}[u]&=-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}u(y)A[v[u]](x,y,s)\right)\\ &\quad+s_{P}\sum_{y\in\mathcal{X}}\hat{q}_{Y}[u](y)\frac{p_{X}(y)}{v[u](y)}\partial f\left(\frac{p_{X}(y)}{v[u](y)}\right)\\ &\quad+s_{P}\Bigg{[}\sum_{y\in\mathcal{X}}\hat{q}_{Y}[u]\left(f\left(\frac{p_{X}(y)}{\hat{q}_{Y}[u](y)}\right)-f\left(\frac{p_{X}(y)}{v[u](y)}\right)\right)\Bigg{]}.\end{split}$ (20)